OpenStudy (anonymous):

What is the pH at the point of neutralization when 25.00 mL of 0.1000 molar acetic acid (Ka = 1.8 x 10-5) reacts with 25.00 mL of 0.1000 molar sodium hydroxide? a) 8.60 b) 8.72 c) 9.76 d) 10.33 e) 11.26

OpenStudy (photon336):

\[pH = pKA + Log[\frac{ [A^{-}] }{ [HA] }]\] Well, the pH at the equivalence point is going to be greater than 7. when titrating with a strong base with a weak acid.

OpenStudy (photon336):

@Cuanchi @aaronq

OpenStudy (anonymous):

so which one is it?

OpenStudy (photon336):

@aaronq what do you think?

OpenStudy (cuanchi):

no, in this case at the equivalence point you dont have the buffer anymore, you only have the salt sodium acetate, then it is a basic salt and you have to calculate the [OH-] \[[OH-]=\sqrt{K _{b} \times C _{s}}\] see example 3 in attached